OpenStudy (omgthatgirl):

What is the equation of a circle with center (3,8) and radius 2? A. (X+3)^2 + (y+8)^2=4 B. (X-3)^2 + (y-8)^2=4 <--- my answer C. (X+3)^2 + (y+8)^2=2 D. (X-3)^2 + (y-8)^2=2

OpenStudy (michele_laino):

hint: we have to apply the definition of a circumference, namely: "the set of all points of the plane such that the distance, from the center, is equal to the radius of such circumference"

OpenStudy (omgthatgirl):

Ok..

OpenStudy (omgthatgirl):

So my answer is wrong

OpenStudy (michele_laino):

it is right!

OpenStudy (omgthatgirl):

K good thanks

OpenStudy (michele_laino):

since, the distance of a point p=(x,y) from the center C=(3,8) is: \[d = \sqrt {{{\left( {x - 3} \right)}^2} + {{\left( {y - 8} \right)}^2}} \] and that distance, has to be equal to the radius of your circumference, namely 2

OpenStudy (michele_laino):

thanks!

OpenStudy (omgthatgirl):

👍